OpenStudy (anonymous):

i am having trouble converting angles into degrees and to radians. how would you work 88 degrees out.

OpenStudy (amistre64):

use the fact that 360 degrees = 2pi radians

OpenStudy (amistre64):

360 = 2pi 1 = 2pi/360 88 = 88*2pi/360

OpenStudy (anonymous):

i see the response, but how to you calculate to get the answer.

OpenStudy (anonymous):

how do you work out 88 degrees.

OpenStudy (amistre64):

i ust worked it out all for you .... not to sure what you are asking about

OpenStudy (anonymous):

Use a calculator to find 88*2pi/360 ... 88/360 = .2444 and 2 pi * .24444 = 1.535 radians

OpenStudy (amistre64):



OpenStudy (amistre64):

88/360 is a part of a circle ... and it gives us a percentage value. 24.44% of 360 degrees converts, or relates, to 24.44% of 2pi

OpenStudy (amistre64):

if you are given radians, then k out of 2pi gives us a percentage to work out on 360 degrees

OpenStudy (anonymous):

how did you come up with 2pi *.24444=1.535

OpenStudy (anonymous):

you have 2 formula: 1/ \(\dfrac{\pi}{180}\) it is used to calculate radian 2/ \( \dfrac{180}{\pi}\) it's used to calculate degree. if you want to calculate radian, use formula 1. For example, you have 88 degree, so convert to radian is 88 * \(\dfrac{\pi}{180}=1.535 ~~radians\) Let say, if you want to convert \(3\pi\) to degree, use formula 2, you will do \(3\pi *\dfrac{180}{\pi}\)= 540 degree. that's it

OpenStudy (anonymous):

pay attention, \(\pi\) is "unit" of radian, so, to calculate radian, it is numerator 180 is "unit" of degree, so , to calculate degree, it is numerator That's the way to memorize and not be confused